A bus leaves a station at 1 P.M., traveling west at an average
rate of 44 mi/h. One hour later a second bus leaves the same station, traveling
east at a rate of 48 mi/h. At what time will the two buses be 274 mi apart?

Answer by checkley75(3666)

44x+48(x-1)=274
44x+48x-48=274
92x=274+48
92x=322
x=322/92
x=3.5 hours the 2 buses will be 274 miles apart
proof
44*3.5+48(3.5-1)=274
154+48*2.5=274
154+120=274
274=274